I don't know where to begin to solve for x and y...please help...
[18 1/24x]=[2x+6 1/4]
[-2/9y 15] [2/3 -5y]

Answer by Jleeblkblt(2)

First find what x and y equal
Then do the first set of brackets [18 1]= [2x+6 1/4] which means 18*1= 18 = 2*x=6 1/4 don't let the slash fool you that means to divide. 1 divided by 4
Move to the second bracket set and do negative 2 divided by 9*y*15 multiply on the second bracket to the first and divide 2 and 3 * negative 5y.
I really hope this answered you question!!!